Two cars leave delhi at the same time ,travelling in opposite directions.if the average speed of one car is 5km/hr more than that of the other and they are 425 km apart at the end of 5hrs,what is the average speed oe each?
Answer by oberobic(2304)

The two cars are moving in opposite directions, so they are separating at the sum of their speeds.
Looking at the fundamental distance equation, d = rt, we know:
d = 425 km
t = 5 hr
So, the combined rate = 425/5 = 85 km/hr
One car goes 5 km/hr faster than the other, which means we can define their speeds as x and x+5.
x + x+5 = 85
2x +5 = 85
2x = 80
x = 40, which is the speed of one car
x+5 = 45, which is the speed of the other car